OpenStudy (anonymous):

What is the slope of the line through the points (-2, -5) and (1, -7)?

OpenStudy (anonymous):

y-y/x-x

OpenStudy (anonymous):

A) m=-2/3 B) m=-3/2 C) m=2/3 D) none of the above

OpenStudy (ashley_f97):

 Use this to help you find your answer

OpenStudy (ashley_f97):

Do you know your x1,x2,y1, and y2?